An Automated Approach to Applying for Employment

George B. Leeman, Jr. IBM Thomas J. Watson Research Center

P. O. Box 704 Yorktown Heights, NY 10598

USA Tel: (914) 784-7354

EMAIL: leeman@watson.ibm.com (Internet)

October 8, 1993 – Version 1.01

Contents

1 Introduction 1 1.1 Rationale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 The Legal Stuff . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 How to Avoid Reading This Report . . . . . . . . . . . . . . . 5

2 A Simple Example of a Tagged File 6

3 Legal Tagged Files 9

4 Four Types of Applicant Files 11 4.1 Use of the IBMRES.STY Templates . . . . . . . . . . . . . . 11 4.2 Use of the NOIBMRES.STY Templates . . . . . . . . . . . . 16 4.3 Use of Tagged Files with Any Formatter or Word Processor . 16 4.4 Use of Raw Text Files . . . . . . . . . . . . . . . . . . . . . . 17

5 Tailoring Resumes to Specific Employers 17

6 User Modifications 18

7 Conclusion 20

1 Introduction

This document describes a piece of software developed at IBM’s Thomas J. Watson Research Center to aid in the employment submission process. The goals of the software and its documentation are:

1. to enable applicants to produce attractive resumes, abstracts, cover letters, and mailing labels with minimal effort;

2. to provide advice on creating effective resumes, including the tailoring of applicant information to specific employers;

3. to lay the foundation for automating the application process by making the initial phases entirely electronic.

This work is based on the notion of a “tagged file” containing data and tags which identify the data; a simple example might be the string “George B. Leeman, Jr.” marked with the tag “\name.” The applicant sends a tagged file to employers who accept such electronic input. However, the applicant may also create paper copies, if they are needed. One may use any formatter or word processor, but special support is provided for the two document preparation systems called TEX and L^{A}TEX, which are extremely popular in universities, industry, and the publishing world. TEX [2] is a sophisticated program capable of producing high-quality typesetting, especially for mathematical text. Its author, Professor Donald E. Knuth of Stanford University, is a remarkable individual whose brilliance, hard work, and gifted writing style have combined to make many long-lasting contributions to the technical community. Leslie Lamport, an internationally well-known computer scientist at Digital Equipment Corporation, has built a collection of TEX commands called L^{A}TEX [3] which capture the document’s structure and hence free the user from dealing with more primitive formatting commands.

1.1 Rationale

Not too long ago the technical publishing world was marked by much waste and duplication of effort. Researchers wrote complicated papers, typically by creating elaborate text files as input to sophisticated formatting programs. The latter produced impressive printed documents which were then sent to journals for publication. If accepted, the journal’s typesetters then retyped the entire paper, essentially duplicating the work of the author! To make matters worse, the galley proofs had to be read by the authors, and errors in complicated mathematical text were inevitable.

Today there are many journals which accept an electronic file of L^{A}TEX commands plus text, and hence the data entry is done only once, by the author. The need to proofread journal galleys is almost eliminated.

Unfortunately there is a procedure in which such waste and duplication still occurs: the application for employment. University students and others seeking jobs produce fancy paper resumes and send them to employers. The employers retype data about the applicants and forward them to possible hiring managers. A large portion of the data entry at the receiving end is duplicating the task already completed by the applicant. Furthermore, the applicant is likely to experience some delay between his request and a reply from the employer; with an electronic system the response could be quicker and sometimes instantaneous.

The main point of this report is to discuss the first set of tools which will make the application process electronic. The applicant creates a tagged file; the form we recommend consists of L^{A}TEX commands and text, but in fact any formatter or word processor can be used. The interface for L^{A}TEX is designed so that the applicant does not have to know this system; he simply fills in the blanks to a template, much like completing a paper form. The file is shipped to employers who can use it for electronic processing, and the user can run L^{A}TEX on the same file to produce attractive paper copies for employers perferring that medium. We also describe other possible tagged files which do not make use of L^{A}TEX. This report gives details on the file creation process and provides tips for creating effective resumes, with the understanding that the recommendations are the opinions of the author and do not necessarily represent a statement of policy by the IBM Corporation.

1.2 The Legal Stuff

Knuth and Lamport made a major decision to give TEX and L^{A}TEX to the world; the source code is available to anyone free of charge. The systems have enjoyed wide exposure, because they are marvellously engineered. However, freedom to examine the code by sympathetic and interested users has also led to corrections and refinements of the systems. We are distributing this software to anyone at no cost, and we hope it will gain wide use in the employment submission process by any employer, not just IBM.

This software resides on an Internet server, called software.watson.ibm- .com, located at the IBM Thomas J. Watson Research Center in Yorktown Heights, NY. A user with access to the Internet may employ the File Transfer Protocol (FTP) to obtain files from such a server. Internet access may require a password; check with the systems administrator at your installation to determine the details. Typically the user logs onto the server remotely with a user name anonymous and requests files from the server. An annotated sample interaction appears in Appendix A. This process of file retrieval is called “anonymous FTP.”

The software described here is essentially a set of new TEX commands (called “macros”—see chapter 20 of [2]). It makes use of a modified form of a L^{A}TEX style file called resume.sty [1] written by Stephen Gildea, with enhancements by Jerry Leichter and this author. The file is available via anonymous FTP from several university servers.

All software on the software.watson.ibm.com server is made available under the following terms and conditions. By downloading the software, you agree to abide by these terms and conditions with respect to this software.

All FTP’able software is copyright International Business Machines Corporation. All rights reserved.

RESEARCH SOFTWARE DISCLAIMER

AS EXPERIMENTAL, RESEARCH SOFTWARE, THIS SET OF TEX MACROS AND DOCUMENTATION ARE PROVIDED FREE OF CHARGE ON AN “AS IS” BASIS WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING BUT NOT LIMITED TO IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. IBM does not warrant that the functions contained in this program will meet the user’s requirements or that the operation of these programs will be uninterrupted or error-free. Acceptance and use of this program constitutes the user’s understanding that he will have no recourse to IBM for any actual or consequential damages, including, but not limited to, lost profits or savings, arising out of the use or inability to use this program, even if the user informs IBM of the possibility of such damages. The user of this program accepts the risk of any harm arising out of the use of this program, or the user shall not attempt to use this program for any purpose.

USER AGREEMENT

BY ACCEPTANCE AND USE OF THIS SET OF TEX MACROS THE USER AGREES TO THE FOLLOWING:

1. These macros are provided for the user’s personal, non-commercial, experimental use, and the user is granted permission to copy these programs to the extent reasonably required for such use.

2. All title, ownership and rights to these macros and any copies remain with IBM, irrespective of the ownership of the media on which the program resides.

3. The user is permitted to create derivative works to this program. However, all copies of the program and its derivative works must contain the IBM copyright notice, the EXPERIMENTAL SOFTWARE DIS-CLAIMER and this USER AGREEMENT.

4. By furnishing this program to the user, IBM does NOT grant either directly or by implication, estoppel, or otherwise any license under any patents, patent applications, trademarks, copyrights, or other rights belonging to IBM or to any third party, except as expressly provided herein.

5. The user understands and agrees that this program and any derivative works are to be used solely for personal, non-commercial, and experimental uses and are not to be otherwise sold, reproduced, leased, transferred, or distributed, or be commercially exploited in any manner.

6. IBM requests that the user supply to IBM a copy of any changes, enhancements, or derivative works which the user may create. The user grants IBM and its subsidiaries an irrevocable, nonexclusive, worldwide and royalty-free license to use, execute, reproduce, display, perform, prepare derivative works based upon, and distribute (INTERNALLY AND EXTERNALLY) copies of any and all such materials and derivative works thereof, and to sublicense others to do any, some, or all of the foregoing (including supporting documentation). Copies of these modifications should be sent to George B. Leeman Jr. at the electronic address watson.ibm.com.

1.3 How to Avoid Reading This Report

In his wonderful user’s guide and reference manual [3], Lamport includes a section with a similar title, and for good reason. Many people prefer to learn a system by playing with it, rather than reading reports. For such individuals we provide sample files with slightly differing interfaces. The files all have the form

/pub/ibmres-tex/file_name.extension,

where the last four characters of the file_name are _c_e, _c_f, _u_e, or _u_f. The letters c and u stand for “commented” and “uncommented,” respectively, and the letters e and f mean “empty” and “filled in,” respectively. The c files have comment lines, identified by a percent sign in column 1, to describe the commands which follow; the u files have few such lines. The e files have no data filled in, while the f files have filled-in examples. Thus all four are skeletons which can be modified, and you may choose the interface you prefer. Using a text editor you then enter appropriate text and give the new file a more sensible name.

There are different sets of files, depending on how the user wants to generate his resumes and other materials.

1. resume in the recommended format for L^{A}TEX. This is the form we recommend, especially for users who have not yet created any resume. The files for this case are named

      /pub/ibmres-tex/res_c_e.ltx, /pub/ibmres-tex/res_c_f.ltx,
      /pub/ibmres-tex/res_u_e.ltx, /pub/ibmres-tex/res_u_f.ltx.

2. resume not in the recommended format, to be used by L^{A}TEX. This form is useful for people who have already created a resume and other materials as a source file for L^{A}TEX and who do not want to retype everything to conform to the format in the preceding case. The files for this second case are named

     /pub/ibmres-tex/nres_c_e.ltx, /pub/ibmres-tex/nres_c_f.ltx,
     /pub/ibmres-tex/nres_u_e.ltx, /pub/ibmres-tex/nres_u_f.ltx.

3. resume not in the recommended format, with tags and ASCII text. This form is useful for people who use other word processors to create their materials. They can fill in the tag data and use the word processor to produce an ASCII text version of their document. The files for this third case are named

     /pub/ibmres-tex/nres_c_e.txt, /pub/ibmres-tex/nres_c_f.txt,
     /pub/ibmres-tex/nres_u_e.txt, /pub/ibmres-tex/nres_u_f.txt.

4. a file with just ASCII text. This form is acceptable, but we do not recommend it. It will have to be examined by a human, whereas the three preceding cases are tagged files which can be processed automatically by software.

If you use the first or second case, while entering text you must beware of ten symbols which have special meaning to TEX and L^{A}TEX: ampersand (&), backslash (\), dollar sign ($), hash mark (#), hat (^), left brace ({), percent sign (%), right brace (}), tilde (~), and underscore (_). You are not likely to want an actual backslash, hat, or tilde in your text, and you may obtain the other seven by typing \ before the symbol, e.g. \& to get an ampersand in the output. Finally, you must figure out how to run L^{A}TEX on the file at your installation if you want paper output.

2 A Simple Example of a Tagged File

In this section we give a simple example which illustrates the use of the ibmres macro package. The first step is to create a text file which looks as follows:

             \documentstyle[12pt,ibmres]{article}
              \begin{document}
                .
                .
                .
              \end{document}

The first line tells L^{A}TEX that you are creating a document in the style of an article, you wish the text to be set in a twelve-point font, and you are using the ibmres macro package. If you prefer a smaller text size, you may substitute 11pt for 12pt; if you omit the string 12pt, you will obtain a document in the default ten-point size. The second and final lines surround the remaining text which is considered the L^{A}TEX document.

We shall call a tag any string which: (a) begins with a backslash \; (b) is immediately followed by a sequence of letters; and (c) optionally ends with a zero. In the example above there are three tags: \documentstyle, \begin, and \end. The case of the letters is significant, so that \end and \End represent different tags. Frequently a tag is followed by a pair of matching braces with a string of text inside; we call that string the argument of the tag. Thus in the preceding examples the string document is the argument of both \begin and \end. An argument may range from the empty string (e.g. \miscellaneous{}) to several lines of text.

Except for lines indicating comments (ignored by L^{A}TEX), which are discussed momentarily, the remainder of the lines in the text file conform to the following format: tag { argument }. Detailed descriptions of the tags appear in subsequent sections, but we give a simple example here.

A text file which includes the 71 lines on pages 21–22 produces the resume, cover letter, and label on pages 23–25. The four initial columns of numbers from 1 to 71 do not actually appear in the input file but are included for easy reference. Lines 1–6 are comments. If L^{A}TEX encounters a %, it ignores any text which remains on that line. As we pointed out in Section 1.3, if you want a % in your text, you must type \%.

The example shows that the tags can be divided into three types. The simplest type, such as \objective (lines 25–26), merely causes specific output (in this case Objective) accompanied by one item (the argument) formatted in an appropriate manner. Note that \name (line 9) and \oneaddressandphone (lines 16–20) are of the same form; the argument of the latter spans all five lines. A second type, such as \education (line 27), implies the occurrence of related sub-tags \when, \place, \location, \degree, \gpa, and \outof (lines 28–33). The third type of tag is illustrated by the string in line 10. The 0 following \preferredtitle means that the tag plus argument will contribute zero to the output; nevertheless, the data will be used by the receiving applicant system. For example, the latter could send back a formal letter in which the argument Mr. Leeman appears in the salutation string Dear Mr. Leeman:. We thus avoid the embarrassing faux pas of using an improper form of address; choosing the right salutation is a task which cannot be done correctly one-hundred percent of the time, even by humans!

L^{A}TEX encourages the user to view the document logically, not physically. The end of line indicator in L^{A}TEX is \\ (double backslash). Accordingly, we shall use the term logical lines to described the lines so separated. Thus the address and phone number on lines 17–19 consist of three logical lines; the equivalent form

                     \oneaddressandphone{
                      2904 Washtenaw
                      Avenue,
                      Apt. 1B\\Ypsilanti,
                      MI 48198\\
                      (313)
                      434-1815
                      }

contains six physical lines for the address and phone number, but still only three logical lines. We shall see examples later where an address and phone number with four logical lines might require five physical lines on output. Our tags are designed so that TEX determines the appropriate physical breaks.

This example illustrates some of the techniques for typing text to be read by L^{A}TEX and TEX.

1. Logical lines are separated by the \\ command. Thus the \\ command in line 18 is essential to force a line break on output; without it, the resume would have appeared at that point as

                Ypsilanti, MI 48198 (313) 434-1815.

The number of physical lines is not important. We prefer a style such as lines 16–20 when the argument has several lines; the closing } is less likely to be omitted. On the other hand, when the argument can be typed on one or two lines, we use a format as in lines 25–26.

2. A blank line such as in lines 55, 59, and 63 causes the end of a paragraph.

3. One can force a blank within a line by the sequence “\ ” (i.e., a backslash followed by a blank), such as in line 69.

As this simple example illustrates, the task of creating effective employment application materials amounts to filling in standard skeletons of tags. One does not have to memorize the special sub-tags associated with tags of the second type above, since they already appear in the skeletons with appropriate comments. When the governing tag (e.g. \education) is not used, then the subordinate tags (e.g. \when…\outof) must be removed as well. These procedures are explored in more detail in the following sections.

We point out that the aforementioned application materials are poorly constructed. Sending a carbon copy to one’s mother is probably a bad idea, and a postscript is questionable in a cover letter for employment, where the tone is more formal. The letter also suffers from the minor defect that it is sent more than twenty years late! The discussion in Section 4.1 points out further deficiencies.

3 Legal Tagged Files

In this section we discuss ten basic tags which should appear in all tagged files. In fact, a tagged file is legal if it contains these tags, plus possibly others. Intuitively the basic tags identify rudimentary information which an employment processing system needs to know about the applicant, such as name, degree, etc. We illustrate the tags with the sample file /pub/ibmres-tex/nres_u_f.txt, a partial listing of which appears on pages 26–27. The tags are as follows.

1. \name (line 8). This tag is usually the first tag, and its argument is your full name.

2. \preferredtitle0 (line 9). As already discussed, it gives the desired string to be used by a responding system in a letter salutation. It produces no output.

3. \school0 (line 10). Enter the school from which you have received or are receiving your most advanced degree. The tag produces no output.

4. \email (line 11). Enter the exact set of symbols which a system could use to respond to you. The form shown is often synonymous with leeman@um.cc.umich.edu. One trap is that the underscore is sometimes present, so one must be careful to type \_, not just _(an example appears in the listing of the file /pub/ibmres-tex/res_u_f.ltx on pages 28–31). If you do not have an electronic mail address, enter

\email{}.

5. \discipline0 (line 12). Enter the discipline in which you received or are receiving your most advanced degree.

6. \highestdegree0 (line 13). Enter the highest degree which you received or are receiving.

7. \degreeyear0 (line 14). Enter the year in which you received or are receiving your highest degree.

8. \workaddressandphone and \homeaddressandphone (examples ap-

pear in lines 13–23 on page 28) or just \oneaddressandphone (lines 15–19). Undergraduates and some graduate students may want to use just one home address via \oneaddressandphone. However, if you have a place of employment, use both a work and home address with

two tags \workaddressandphone and \homeaddressandphone. A fre-

quently asked question is whether a home address refers to the living quarters at a university or a more permanent site, such as the home of a student’s parents. If you will be at a campus residence six weeks after you send out your resumes, then use the campus residence; otherwise use a more permanent address. Separate logical lines by \\; the system will figure out proper physical line breaks.

9. \immigrationstatus and \typeofpositiondesired (lines 20–21).

These extremely important entries are often not found at all in resumes! They are always key elements which a potential hiring manager must know and should therefore appear early.

(a) \immigrationstatus. Typical entries are US citizen, US permanent resident, Student Visa. Some students prefer to indicate the type of visa (e.g. F1, H1, J1), since they carry different implications for employment. This entry is also important for the hiring manager to know. For example, employers offering work involving US government contracts may be forced to consider only US citizens, because those in the other categories may be unable to obtain a government security clearance.

(b) \typeofpositiondesired. Entries here are Permanent, Summer (indicate the year—an omission in the example of Section 2), and Temporary; an example of the latter is a co-op position for a semester.

4 Four Types of Applicant Files

In this section we discuss four possible formats for electronic files. The first three are legal tagged files which include the basic tags presented in the preceding section. The final format is just raw, unstructured text made up of ASCII characters.

4.1 Use of the IBMRES.STY Templates

If you have not created a resume and associated materials, we recommend using the ibmres.sty approach. You fill in one of the templates listed in case 1 of Section 1.3. If you want paper versions of your document, you ensure that the file /pub/ibmres-tex/ibmres.sty is available, and you run L^{A}TEX on your file. We provide detailed descriptions of the various tags available, and we give tips on producing effective resumes. We illustrate the tags with a more complicated example of a PhD student resume. The only tags which are required are the basic tags mentioned in the preceding section. However, many of the others which are discussed here will be useful. The input is the file /pub/ibmres-tex/res_u_f.ltx (pages 28–31), and the output is on pages 32–36. In this discussion two points will be emphasized:

1. Both the choice of items in a resume and their order on the page are important. You should express your qualifications in a terse but convincing style, and should arrange items so that the potential hiring manager will see key points early.

2. You may want to create several resumes, tailored to specific employers. However, begin with a general version, which can then be slightly modified by adding or removing various tags.

We now discuss one-by-one the remaining tags which have not already been mentioned in the preceding section. Although in the resume output the left margin is really the blank 1′′x11′′ rectangle extending from the left side of the page, we shall instead call the left margin the rectangle which starts at the beginning of Objective and ends at the beginning of “To obtain a research …” in our example.

1. \presentstatus (line 28) and \availabledate (line 31). These extremely important entries are often not found at all in resumes! They are always key elements which a potential hiring manager must know and should therefore appear early.

(a) \presentstatus. Your present status might be BS student, MS student, PhD student, Postdoc student, faculty, employed, etc. This category often immediately determines whether you are a candidate for an available position.

(b) \availabledate. This item is a possible starting date. Your best strategy is to put the earliest possible date which you can tolerate, provided that all obligations can be fulfilled. A counterexample would be a date July 1 and an education item saying PhD expected July 1, where the latter date is unrealistically optimistic.

2. \employerspecifictitle, \employerspecificdata (lines 32–33).

These specialized tags are only used in a minority of resumes, and hence they are discussed in the next section.

3. \objective (lines 34–35). The objective must be clear and concise. BS students tend to have more general objectives (e.g. “a position in software engineering”), whereas PhDs are more specific (see example in lines 34–35). Keep in mind that the goal is to write a resume in which the items that follow illustrate that you are well-qualified to pursue the objective.

4. \education (line 36). This tag spawns one or more sextets consisting

of the tags \when, \place, \location, \degree, \gpa, \outof.

(a) \when (line 37). When the degree was received or when it is expected.

(b) \place (line 38). The university.

(c) \location (line 39). Its location; in the US, city and state suffice.

(d) \degree (lines 40–41). Include the type (e.g. BA, MA, BS, MS, PhD, etc.) and the field of major study. When appropriate, include a thesis advisor and title. The thesis advisor may also be listed in the References section, which is a desirable redundancy.

(e) \gpa (line 42). Usually a decimal number, but there may be other cases (see line 54).

(f) \outof (line 43). The highest possible GPA. This entry is important and sometimes omitted. Many graduate schools use a maximum of 9.000. If you type \gpa{7.362}\outof{} instead of \gpa{7.362}\outof{8.000}, the reader may assume the omitted field was 9.000 and consider you a B student instead of an A student.

5. \employment (line 56). If you have employment experience, use this tag, which implies one of more quartets of the tags \when, \place,

\location, \text.

(a) \when (line 57). A range of dates for the job.

(b) \place (line 58). Name of employer.

(c) \location (line 59). Location; in the US, city and state suffice.

(d) \text (lines 60–61). Text giving a brief description of the work. If the work is related to the Objective, choose wording which makes the connection clear.

6. \skills (line 66). This is an optional tag which is a catch-all for capabilities which do not appear anywhere else. It is followed by one or more \text tags to describe the items.

(a) \text (line 67). As an example, the ability to write programs in a specific programming language may be an important adjunct for carrying out the objective. For some disciplines the nature of your degree and education might be sufficient, and the \skills and \text tags could be omitted.

7. \honors (line 68). If you have honors to report, use this tag followed by one or more \text tags.

(a) \text (line 69). Appellations such as magna cum laude and summa cum laude are often listed as part of the degree in the Education section, but honors buried there are apt to be missed. Special employer-specific honors are best handled in a different manner (see next section).

8. \publications (line 72). If you have publications, use this tag followed by one or more \text tags.

(a) \text (line 73). List title, authors, and when and where the paper appeared or will appear. Any format seen in bibliographies is acceptable; the form shown here with no font changes is adequate. It is important to observe that the number of publications may depend not only on the student, but also the discipline and the advisor. A pure mathematics PhD writing a first-rate thesis might derive one or two publications from it in serious journals; an engineering student with results of immediate applicability might be encouraged to write up portions and present them at a conference, resulting in a larger number of publications.

9. \miscellaneous (line 85). This is an entry which can be omitted, and its presence can affect the tone of the resume. In this writer’s opinion, items such as “enjoys tennis, classical music, and house building” are less effective than entries such as significant volunteer work or positions of authority and/or responsibility.

10. \references (line 87). This tag is obligatory and gives rise to one

or more pairs \referencename and \referenceemail. Writing some-

thing like “provided upon request” (as was done in the example of Section 2) is a mistake. Any potentially interested employer will ask for references, so such an entry only promotes delay. Furthermore, a list of names has two additional benefits. First, the list may suggest a breadth which the Objective requires. For instance, in the example of Section 2 the applicant could have listed both a physicist and a computer scientist, or both a mathematician and a computer scientist among the references. Second, the manager might recognize one of the names, especially if the reference is well-known.

(a) \referencename (line 88). Include the full name, title, location, and phone number.

(b) \referenceemail (line 90). Include the exact set of characters in the electronic address; make the argument null if the reference has no such address. Note that an automated system could request reference letters automatically, and the reference could forward them in electronic form.

11. \coverletter (lines 97–102). The impressive business letter is similar to, but not as sophisticated as, the sample format written by Knuth [2, pp. 403–408]. As already discussed, this tag’s argument consists of at least two logical lines separated by \\ commands. The first line appears at the top of the letterhead in a large font; the remaining lines are joined with a comma and appear underneath the first line. If the argument is null, the next eight tags have no effect, even if their arguments are not null.

(a) \rightlines (line 103–107). The logical lines of the argument appear at the right-hand side of the business letter, separated from the preceding entries by a long rule. Typical choices here are the sender’s name and title, electronic address, and phone number.

(b) \recipient (lines 108–113). Include the addressee’s name and/or title and address.

(c) \letterbody (lines 114–126). Separate paragraphs with blank lines (lines 116, 120, 124). This writer prefers extra space between paragraphs, which was obtained via the \bigskip TEX command (lines 117, 121, 125). It is useful to include any information here which does not conveniently fit in the resume. In particular, providing a more detailed account of your availability can save the employer time in trying to contact you. The date of the letter is generated automatically.

(d) \closing (lines 127–130). Include a suitable closing and your name, separated by \\.

(e) \cc (line 131). The carbon copy list has limited use in a cover letter. It can be omitted or equivalently used with a null argument. Members of a cc: list would be separated by \\.

(f) \encl (line 132). It is a good idea to list as enclosures all materials accompanying the letter. Separate items in the list with \\.

(g) \ps (line 133). As mentioned previously, use of the postscript tag can affect the tone of the letter. It may be omitted or equivalently used with null argument.

(h) \letterlabel (line 134). This tag produces a mailing label. The return address is generated from the letterhead, and the remainder of the label is the \recipient argument.

12. \abstracts (line 135). If you want to include some abstracts, follow this tag by one or more \text tags. A good strategy is to fit them on one page. Large numbers of pages will inundate the employer. A second strategy is to supply no abstracts when the resume is first sent, but to be ready with the additional page as a follow-up weapon.

(a) \text (line 136–146). It would take us too far afield to explain all the typed input in these examples. However, they show how easily mathematical formulas can be obtained via L^{A}TEX and TEX.

4.2 Use of the NOIBMRES.STY Templates

Although the ibmres.sty package in the preceding section is the one we recommend most, there are cases in which it may not be the best choice. If you are already using L^{A}TEX for resumes and related materials, you may not want to retype everything to cater to the ibmres format. In that case make sure the file noibmres.sty (available via anonymous FTP from software.watson.ibm.com) is part of your system, and create a file which includes the basic tags plus your own L^{A}TEX source file. An example appears in the listing on page 37, where your own L^{A}TEX commands and text would appear after line 24. There is only one string attached—your L^{A}TEX source must be self-contained, i.e., not reference any special style or macro files. You may of course include any needed macro definitions between lines 8 and 9. The noibmres.sty file simply defines each tag to produce no text. If your text after line 24 happens to have a basic tag such as \name{...}, you may omit that tag (line 10 in the example) at the beginning of the file. If you do not, the receiving software will warn you that two identical basic tags were detected, and it will use the data associated with the first occurrence. In other words, you may redefine any of the basic tags in your own L^{A}TEX source, and they may appear in any order. Any which do appear in your text may be omitted at the beginning of the file (lines 10–23 in our example).

4.3 Use of Tagged Files with Any Formatter or Word Processor

We strongly recommend the use of L^{A}TEX for creating resumes and related items, because it is ubiquitous and produces beautiful results. However, if you have already created your materials with some other formatter or word processor, you may prefer not to retype everything to conform to L^{A}TEX. If the program you used has the capability to produce a raw ASCII text version of your documents, you may fill in a template with the basic tags and append that version at the end. An example appears on pages 26–27. With this approach you avoid duplicate typing but still derive the benefits of submitting a structured file which can be quickly processed by receiving software.

4.4 Use of Raw Text Files

The final possibility—in our opinion the poorest choice of the four cases—is to submit an unstructured file with just ASCII text. We discourage this approach, because with minimal effort you can use the preceding alternative (at the beginning of your unstructured file, include the basic tags with appropriate data). Nevertheless, we feel that even this fourth choice is better than sending paper copies. Electronic submission of any kind is fast, and your materials can be forwarded in the same form to multiple locations quickly.

5 Tailoring Resumes to Specific Employers

In this section we show how to use the ibmres.sty package to tailor your general resume to a specific employer. The most common reason is that you have either worked for that employer in the past or you have achievements which might be of greater interest to that employer. The best location for such material is right above the Objective, so that the reader sees it immediately. Two tags are provided, which we will illustrate with examples.

1. \employerspecifictitle. Indicate a title for the item. Two examples for the IBM corporation are:

           \employerspecifictitle{IBM Fellowship Holder:}
           \employerspecifictitle{IBM Work Experience:}

2. \employerspecificdata. Indicate data which explains the title. For the two examples above one might use

             \employerspecificdata{dates of fellowship}

and

        \employerspecifictitle{location, date, manager name}

Note there is no problem with redundancy between these items and the Employment section. The manager name would be of great interest to another IBM manager, who could then get an inside opinion about the applicant. It would make less sense to list that name on a general resume sent to other employers.

Another example is any honor which is special to the employer. In the case of IBM, if you won a prize in an IBM-sponsored contest, the award would have greater weight in this early position on the resume, especially if accompanied by the name of someone in IBM involved with the contest.

These two tags occur in pairs and are formatted like the \presentstatus tag and its argument. The argument to \employerspecifictitle is in bold face, and the argument of \employerspecificdata appears immediately afterward.

6 User Modifications

In this section we show how one can make simple changes to the ibmres.sty package. Occasionally users want to add a tag which might not be present. The easiest way is to plunge into the file /pub/ibmres-tex/ibmres.sty to see how the existing tags are formed. That task is made less formidable by including some examples here.

It is important that you add your changes in the file you will be sending, not (for example) /pub/ibmres-tex/ibmres.sty. The right place to add new definitions is between the \documentstyle line and the \begin{document} line. The recipient can then still run L^{A}TEX on your file with his own version of /pub/ibmres-tex/ibmres.sty.

Suppose you are applying for an academic position and feel that a list of advanced graduate courses you have taken would enhance your resume. In terms of formatting this task is no different than designing the \skills and \honors tags, since you want a bold face entry in the left margin with the possibility of one or more items to the right. The latter piece can already be done with the \text tag. To attack the former piece, we find in the file /pub/ibmres-tex/ibmres.sty the lines

          \def\skills#1{\sectiontitle{Skills}#1\par}
          \def\honors#1{\sectiontitle{Honors}#1\par}

where \def is a TEX command defining the control sequence (\skills and \honors in these cases) which immediately follows. The big advantage of examining two instances is that you see what portions change and what portions remain the same. Even if you know little about TEX, you can guess that the configuration

  \def\courses#1{\sectiontitle{Graduate Courses Taken}#1\par}

will yield Graduate Courses Taken in the left margin, and thus the sequence

   \courses{}
   \text{Several Complex Variables, Distribution Theory, ...}

in our input file achieves the desired effect. You add the new \def line immediately after the \documentstyle line.

If you want two lists—one for courses taken and one for courses taught— you may define two new tags \taken and \taught by

       \def\taken#1{\sectiontitle{Courses Taken}#1\par}
      \def\taught#1{\sectiontitle{Courses Taught}#1\par}

which you add at the top of the file after the \documentstyle line.

Similarly, if you want a new set of tags in which the heading appears horizontally above the data and part of the data falls in the left margin, you can mimic the \education and \when tags to get the desired effect.

We mention one more physical characteristic which you might want to change: the width of the left margin. The last two lines in /pub/ibmrestex/ibmres.sty

       \settowidth{\leftmarginwidth}{\bf Miscellaneous}
        \addtolength{\leftmarginwidth}{1 em}

set an internal dimension called \leftmarginwidth to the precise horizontal width of the word Miscellaneous in boldface, then add a unit of 1 em to that dimension. If you are not using the \miscellaneous tag and decide that Publications is the widest title in the left margin, you may insert the lines

        \settowidth{\leftmarginwidth}{\bf Publications}
        \addtolength{\leftmarginwidth}{1 em}

after the \documentstyle line.

The entire listing of the /pub/ibmres-tex/ibmres.sty file is found in Appendix B. We exhort the reader to become more familiar with L^{A}TEX and TEX and to suggest improvements to this file.

7 Conclusion

This report has discussed an approach to simplify the initial phases of the employment application process and make them electronic. Several other tools and documentation would be useful. Examples include parsers which check the created input file for accuracy; systems which immediately acknowledge the applicant’s submission and modify and distribute the incoming data appropriately; and further discussion about and improvements to the macros. Work on some of these pieces is nearing completion.

The author would like to thank Alan D. Wilson, for supporting this work; Carol A. Johnson, Michael Sampogna, and John G. Oxton, for insightful discussions about recruiting operations; Lawrence A. Gallo and Ryuichi Sagara, for building a first-rate recruiting system which will accelerate the implementation of the ideas in this report; and Michael S. Mosley and Blaney Harper, who suggested a number of manuscript improvements.

Listing of /pub/ibmres-tex/simpres.ltx File

001 %                                                             10/08/93 %
002 % IBM simple resume file, /pub/ibmres-tex/simpres.ltx                   %
003 % George B. Leeman, Jr., leeman@watson.ibm.com.                         %
004 % Please read /pub/README.FTP on software.watson.ibm.com before using   %
005 % this file.  Run LaTeX on this file to produce a simple resume, cover  %
006 % letter, and mailing label.                                           %
007 \documentstyle[12pt,ibmres]{article}
008 \begin{document}
009 \name{George B. Leeman, Jr.}
010 \preferredtitle0{Mr. Leeman}
011 \school0{The University of Michigan}
012 \email{leeman at um.cc.umich.edu}
013 \discipline0{mathematics}
014 \highestdegree0{PhD}
015 \degreeyear0{1972}
016 \oneaddressandphone{
017 2904 Washtenaw Avenue, Apt. 1B\\
018 Ypsilanti, MI 48198\\
019 (313) 434-1815
020 }
021 \presentstatus{BS Student}
022 \immigrationstatus{US citizen}
023 \typeofpositiondesired{Summer}
024 \availabledate{June, 1973}
025 \objective{To obtain a position in numerical analysis and/or scientific
026 programming.}
027 \education{}
028 \when{May 1973}
029 \place{The University of Michigan}
030 \location{Ann Arbor, MI}
031 \degree{BA in Mathematics.}
032 \gpa{3.603}
033 \outof{4.000}
034 \references{}
035 \referencename{Provided upon request.}
036 \referenceemail{}
037 \coverletter{
038 The University of Michigan\\
039 Department of Mathematics\\
040 Ann Arbor, MI 48109-2122
041 }

22 LISTING OF /PUB/IBMRES-TEX/SIMPRES.LTX FILE

042 \rightlines{
043 George B. Leeman, Jr.\\
044 leeman at um.cc.umich.edu\\
045 (313) 764-8504
046 }
047 \recipient{
048 Manager, Mid-Hudson Valley Recruiting\\
049 Employment Solutions Corporation\\
050 P. O. Box 390\\
051 Poughkeepsie, New York 12602
052 }
053 \letterbody{
054 Dear Sir:
055
056 I would like to apply for a summer, 1973 position within the IBM
057 development divisions in the Mid-Hudson Valley region.  A resume is
058 included.
059
060 I can be reached at the number shown above every afternoon from 1:00 PM
061 to 5:00 PM.  I answer electronic mail throughout each day, including
062 weekends.
063
064 Thank you for your consideration.
065 }
066 \closing{Sincerely yours,\\George B. Leeman, Jr.}
067 \cc{my mother}
068 \encl{resume}
069 \ps{P.\ S.\ Thank you for making it possible to send resumes electronically!}
070 \letterlabel{}
071 \end{document}

George B. Leeman, Jr.

2904 Washtenaw Avenue, Apt. 1B Ypsilanti, MI 48198

(313) 434-1815

leeman at um.cc.umich.edu

Present Status: BS Student Type of Position Desired: Summer Immigration Status: US citizen Available Date: June, 1973

Objective To obtain a position in numerical analysis and/or scientific programming.

Education

May 1973  The University of Michigan  Ann Arbor, MI

BA in Mathematics. GPA: 3.603 out of 4.000.

References Provided upon request.

The University of Michigan

Department of Mathematics, Ann Arbor, MI 48109-2122 George B. Leeman, Jr. leeman at um.cc.umich.edu (313) 764-8504

June 12, 1997

Manager, Mid-Hudson Valley Recruiting Employment Solutions Corporation P. O. Box 390 Poughkeepsie, New York 12602

Dear Sir: I would like to apply for a summer, 1973 position within the IBM development divisions in the Mid-Hudson Valley region. A resume is included. I can be reached at the number shown above every afternoon from 1:00 PM to 5:00 PM. I answer electronic mail throughout each day, including weekends. Thank you for your consideration.

Sincerely yours,

George B. Leeman, Jr.

cc: my mother encl: resume P. S. Thank you for making it possible to send resumes electronically!

George B. Leeman, Jr. The University of Michigan Department of Mathematics Ann Arbor, MI 48109-2122

Manager, Mid-Hudson Valley Recruiting Employment Solutions Corporation P. O. Box 390 Poughkeepsie, New York 12602 26 PARTIAL LISTING OF /PUB/IBMRES-TEX/NRES_U_F.TXT FILE

Partial Listing of /pub/ibmres-tex/nres_u_f.txt File

001 %                                                             10/08/93 %
002 % NON-latex RESume file, Uncommented, Filled in,                        %
003 % /pub/ibmres-tex/nres_u_f.txt                                         %
004 % George B. Leeman, Jr., leeman@watson.ibm.com.                         %
005 % Please read /pub/README.FTP on software.watson.ibm.com before using   %
006 % this file.  Change the text in the arguments, then add your own text  %
007 % at the end.                                                          %
008 \name{George B. Leeman, Jr.}
009 \preferredtitle0{Mr. Leeman}
010 \school0{The University of Michigan}
011 \email{leeman at um.cc.umich.edu}
012 \discipline0{mathematics}
013 \highestdegree0{PhD}
014 \degreeyear0{1972}
015 \oneaddressandphone{
016 2904 Washtenaw Avenue, Apt. 1B\\
017 Ypsilanti, MI 48198\\
018 (313) 434-1815
019 }
020 \immigrationstatus{US citizen}
021 \typeofpositiondesired{Permanent}
022                             George B. Leeman, Jr.
023
024 Work Address                                   Home Address
025 The University of Michigan                      2904 Washtenaw Avenue, Apt. 1B
026 Department of Electrical Engineering            Ypsilanti, MI 48198
027  and Computer Science                           (313) 434-1815
028 Ann Arbor, MI 48109-2122
029 (313) 764--8504
030 leeman@um.cc.umich.edu
031
032 Present Status: Postdoc Student                 Immigration Status: US citizen
033
034 Type of Position Desired: Permanent             Available Date: June, 1973
035
036 Objective      To obtain a research or development position in the areas of
037                stochastic analysis or program correctness.
  .
  .
  .
100 The University of Michigan

101 Department of Electrical Engineering and Computer Science, 1301 Beal Avenue,
102 Ann Arbor, MI 48109-2122
103 ------------------------------------------------------------------------------
104 George B. Leeman, Jr.
105 leeman at um.cc.umich.edu
106 (313) 764-8504
107                                                            September 21, 1993
108 Manager, PhD Recruiting
109
110 IBM Thomas J. Watson Research Center
111 P. O. Box 218
112 Yorktown Heights, New York 10598
113
114 Dear Sir:
115
116 I would like to apply for a position in the research and development divisions
117 of your corporation. I have included a resume and if necessary can send a few
118 abstracts from some of my published papers.
119
120 I can be reached at the number shown above every afternoon from 1:00 P.M. to
121 5:00 P.M. I answer electronic mail throughout each day, including weekends.
122
123 Thank you for your consideration.
124
125 Sincerely yours,
126
127 George B. Leeman, Jr.
128 encl: resume

28 LISTING OF /PUB/IBMRES-TEX/RES_U_F.LTX FILE

Listing of /pub/ibmres-tex/res_u_f.ltx File

001 %                                                             10/08/93 %
002 % ibm RESume file, Uncommented, Filled in, /pub/ibmres-tex/res_u_f.ltx  %
003 % George B. Leeman, Jr., leeman@watson.ibm.com.                         %
004 % Please read /pub/README.FTP on software.watson.ibm.com before using   %
005 % this file.  Change the text in the arguments, then run LaTeX on this  %
006 % file to produce resume, cover letter,  mailing label, and selected    %
007 % abstracts.                                                           %
008 \documentstyle[12pt,ibmres]{article}
009 \begin{document}
010 \name{George B. Leeman, Jr.}
011 \preferredtitle0{Mr. Leeman}
012 \school0{The University of Michigan}
013 \email{leeman@um.cc.umich.edu}
014 \discipline0{mathematics}
015 \highestdegree0{PhD}
016 \degreeyear0{1972}
017 \workaddressandphone{
018 The University of Michigan\\
019 Department of Electrical Engineering and Computer Science\\
020 Ann Arbor, MI 48109-2122\\
021 (313) 764--8504
022 }
023 \homeaddressandphone{
024 2904 Washtenaw Avenue, Apt. 1B\\
025 Ypsilanti, MI 48198\\
026 (313) 434-1815
027 }
028 \presentstatus{Postdoc Student}
029 \immigrationstatus{US citizen}
030 \typeofpositiondesired{Permanent}
031 \availabledate{June, 1973}
032 \employerspecifictitle{}
033 \employerspecificdata{}
034 \objective{To obtain a research or development position in the areas of
035 stochastic analysis or program correctness.}
036 \education{}
037 \when{Aug. 1972}
038 \place{The University of Michigan}
039 \location{Ann Arbor, MI}
040 \degree{PhD in Mathematics; advisor: Peter L. Duren; thesis title:
041 The constrained coefficient problem for typically real functions.}

042 \gpa{8.217}
043 \outof{9.000}
044 \when{May 1969}
045 \place{The University of Michigan}
046 \location{Ann Arbor, MI}
047 \degree{MA in Mathematics}
048 \gpa{8.041}
049 \outof{9.000}
050 \when{June 1968}
051 \place{Yale University}
052 \location{New Haven, CT}
053 \degree{BA in Mathematics}
054 \gpa{abolished 11/30/67; class percentile: 97.}
055 \outof{}
056 \employment{}
057 \when{summers, 1970, 1971}
058 \place{The University of Michigan}
059 \location{Ann Arbor, MI}
060 \text{Instructor, teaching undergraduate introductory and intermediate
061 calculus courses.}
062 \when{summers, 1965--1969}
063 \place{Perkin-Elmer Corporation}
064 \location{Norwalk, CT}
065 \text{Scientific programming for problems in engineering and physics.}
066 \skills{}
067 \text{Programming in FORTRAN, IBM 360 Assembler, SNOBOL4, LISP, C.}
068 \honors{}
069 \text{National Science Foundation Traineeship, September, 1968 to August, 1972.}
070 \text{BA Magna Cum Laude, honors with exceptional distinction, 1968.}
071 \text{Election to Phi Beta Kappa, November, 1967.}
072 \publications{}
073 \text{The seventh coefficient of odd symmetric univalent functions,
074 G. B. Leeman, Jr.,
075 to appear in Duke Mathematical Journal, vol. 43, no. 2, June,
076 1973.}
077 \text{A new proof for an inequality of Jenkins,
078 G. B. Leeman, Jr.,
079 Proceedings of the American Mathematical Society, vol. 54, Jan.
080 1973, 114--116.}
081 \text{The constrained coefficient problem for typically real functions,
082 G. B. Leeman, Jr.,
083 Transactions of the American Mathematical Society, vol. 186,
084 Dec. 1972, 177--189.}
085 \miscellaneous{Member of Board of Directors, Ridgefield Symphony Orchestra,

30 LISTING OF /PUB/IBMRES-TEX/RES_U_F.LTX FILE

086 Ridgefield, CT.}
087 \references{}
088 \referencename{Peter L. Duren, Professor, The University of Michigan,
089 (313) 764-0202.}
090 \referenceemail{duren@um.cc.umich.edu}
091 \referencename{Bernard A. Galler, Professor, The University of Michigan,
092 (313) 764-5832.}
093 \referenceemail{bernard\_a.\_galler@um.cc.umich.edu}
094 \referencename{Maxwell O. Reade, Professor, The University of Michigan,
095 (313) 764-7227.}
096 \referenceemail{}
097 \coverletter{
098 The University of Michigan\\
099 Department of Electrical Engineering and Computer Science\\
100 1301 Beal Avenue\\
101 Ann Arbor, MI 48109-2122
102 }
103 \rightlines{
104 George B. Leeman, Jr.\\
105 leeman at um.cc.umich.edu\\
106 (313) 764-8504
107 }
108 \recipient{
109 Manager, PhD Recruiting\\
110 IBM Thomas J. Watson Research Center\\
111 P. O. Box 218\\
112 Yorktown Heights,  New York 10598
113 }
114 \letterbody{
115 Dear Sir:
116
117 \bigskip I would like to apply for a position in the research and development
118 divisions of your corporation.  I have included a resume and a few abstracts
119 from some of my published papers.
120
121 \bigskip I can be reached at the number shown above every afternoon from
122 1:00 P.M. to 5:00 P.M.  I answer electronic mail throughout each day, including
123 weekends.
124
125 \bigskip Thank you for your consideration.
126 }
127 \closing{
128 Sincerely yours,\\
129 George B. Leeman, Jr.

130 }
131 \cc{}
132 \encl{resume\\selected abstracts}
133 \ps{}
134 \letterlabel{}
135 \abstracts{}
136 \text{
137 {\em The seventh coefficient of odd symmetric univalent functions, by
138 G. B. Leeman, Jr.}\vskip 3ex
139 Let $S_{odd}$ be the collection of all functions $f(z) = z +
140 c_3z^3 + c_5z^5 + c_7z^7 + \cdots$ odd, analytic, and one-to-one in the unit
141 disk.  In 1933 Fekete and Szeg\"o showed that for all $f$ in $S_{odd}$, $|c_5|
142 \leq 1/2 + e^{-2/3}$, but no sharp bounds have been found since that time, even
143 for the subclass of $S_{odd}$ with real coefficients.  In this paper we find the
144 sharp bound $|c_7|\leq 1090/1083$ for this subclass, and we identify all
145 extremal functions.\vskip 7 ex
146 }
147 \text{
148 {\em A new proof for an inequality of Jenkins, by G. B. Leeman, Jr.}\vskip 3 ex
149 A new proof of Jenkins' inequality
150 $${\rm Re}(e^{2i\theta}a_3 - e^{2i\theta} a_2^2 - \tau e^{i\theta}a_2) \leq 1 +
151 {3\over8} \tau^2 - {1\over4}\tau^2 \log ({\tau\over4}),\ \ 0 \leq \tau \leq 4,$$
152 for univalent functions $f(z) = z + \sum_{n=2}^\infty a_n z^n$ is
153 presented.\vskip 7 ex
154 }
155 \text{
156 {\em The constrained coefficient problem for typically real functions, by
157 G. B. Leeman, Jr.}\vskip 3 ex
158 Let $-2 \leq c \leq 2$.  In this paper we find the precise upper
159 and lower bounds on the $n$th Taylor coefficient $a_n$ of functions $f(z) = z +
160 c z^2 + \sum_{k=3}^\infty a_k z^k$ typically real in the unit disk for
161 $n=3,4,\cdots \ .$  In addition all the extremal functions are identified.
162 }
163 \end{document}

George B. Leeman, Jr.

Work Address The University of Michigan Department of Electrical Engineering

and Computer Science Ann Arbor, MI 48109-2122

(313) 764–8504 leeman@um.cc.umich.edu Home Address 2904 Washtenaw Avenue, Apt. 1B Ypsilanti, MI 48198

(313) 434-1815

Present Status: Postdoc Student Type of Position Desired: Permanent Immigration Status: US citizen Available Date: June, 1973

Objective To obtain a research or development position in the areas of stochastic analysis or program correctness.

Education

Aug. 1972  The University of Michigan  Ann Arbor, MI

PhD in Mathematics; advisor: Peter L. Duren; thesis title: The constrained coefficient problem for typically real functions. GPA: 8.217 out of 9.000.

May 1969  The University of Michigan  Ann Arbor, MI

MA in Mathematics GPA: 8.041 out of 9.000.

June 1968  Yale University  New Haven, CT

BA in Mathematics GPA: abolished 11/30/67; class percentile: 97.

Employment

{summers, 1970, \atop 1971} The University of Michigan Ann Arbor, MI Instructor, teaching undergraduate introductory and intermediate calculus courses.

{summers, \atop 1965–1969}  Perkin-Elmer Corporation  Norwalk, CT

Scientific programming for problems in engineering and physics.

Skills Programming in FORTRAN, IBM 360 Assembler, SNOBOL4, LISP, C.

Honors National Science Foundation Traineeship, September, 1968 to August, 1972.

BA Magna Cum Laude, honors with exceptional distinction, 1968.

Election to Phi Beta Kappa, November, 1967.

Publications The seventh coefficient of odd symmetric univalent functions, G. B. Leeman, Jr., to appear in Duke Mathematical Journal, vol. 43, no. 2, June, 1973.

A new proof for an inequality of Jenkins, G. B. Leeman, Jr., Proceedings of the American Mathematical Society, vol. 54, Jan. 1973, 114–116.

The constrained coefficient problem for typically real functions, G. B. Leeman, Jr., Transactions of the American Mathematical Society, vol. 186, Dec. 1972, 177–189.

Miscellaneous Member of Board of Directors, Ridgefield Symphony Orchestra, Ridgefield, CT.

References Peter L. Duren, Professor, The University of Michigan, (313) 764-0202. Email: duren@um.cc.umich.edu

Bernard A. Galler, Professor, The University of Michigan, (313) 764-5832. Email: bernard_a._galler@um.cc.umich.edu

Maxwell O. Reade, Professor, The University of Michigan, (313) 764-7227.

The University of Michigan

Department of Electrical Engineering and Computer Science, 1301 Beal Avenue, Ann Arbor, MI 48109-2122 George B. Leeman, Jr. leeman at um.cc.umich.edu (313) 764-8504

June 12, 1997

Manager, PhD Recruiting IBM Thomas J. Watson Research Center P. O. Box 218 Yorktown Heights, New York 10598

Dear Sir:

I would like to apply for a position in the research and development divisions of your corporation. I have included a resume and a few abstracts from some of my published papers.

I can be reached at the number shown above every afternoon from 1:00 P.M. to 5:00 P.M. I answer electronic mail throughout each day, including weekends.

Thank you for your consideration.

Sincerely yours,

George B. Leeman, Jr.

encl: resume selected abstracts

George B. Leeman, Jr. The University of Michigan Department of Electrical Engineering and Computer Science 1301 Beal Avenue Ann Arbor, MI 48109-2122

Manager, PhD Recruiting

IBM Thomas J. Watson Research Center P. O. Box 218 Yorktown Heights, New York 10598

Selected Abstracts

The seventh coefficient of odd symmetric univalent functions, by G. B. Leeman, Jr.

Let S_{odd} be the collection of all functions f(z) = z +c_{3}z^{3} +c_{5}z^{5} +c_{7}z^{7} +⋯ odd, analytic, and one-to-one in the unit disk. In 1933 Fekete and Szegö showed that for all f in S_{odd}, |c_{5}| ≤ 1/2 + e^{−2/3}, but no sharp bounds have been found since that time, even for the subclass of S_{odd} with real coefficients. In this paper we find the sharp bound |c_{7}| ≤ 1090/1083 for this subclass, and we identify all extremal functions.

A new proof for an inequality of Jenkins, by G. B. Leeman, Jr.

A new proof of Jenkins’ inequality

Re(e^{2iθ}a_{3} − e^{2iθ}a_{2}^{2} − τe^{iθ}a_{2}) ≤ 1 + \frac{3}{8}τ^{2} − \frac{1}{4}τ^{2} log(\frac{τ}{4}), 0 ≤ τ ≤ 4,

for univalent functions f(z) = z + ∑_{n=2}^{∞} a_{n}z^{n} is presented.

The constrained coefficient problem for typically real functions, by G. B. Leeman, Jr.

Let −2 ≤ c ≤ 2. In this paper we find the precise upper and lower bounds on the nth Taylor coefficient a_{n} of functions f(z) = z + cz^{2} + ∑_{k=3}^{∞} a_{k}z^{k} typically real in the unit disk for n = 3, 4, · · · . In addition all the extremal functions are identified.

Listing of /pub/ibmres-tex/nres_u_f.ltx File

001 %                                                             10/08/93 %
002 % Non-ibm RESume file, Uncommented, Filled in,                          %
003 % /pub/ibmres-tex/nres_u_f.ltx                                         %
004 % George B. Leeman, Jr., leeman@watson.ibm.com.                         %
005 % Please read /pub/README.FTP on software.watson.ibm.com before using   %
006 % this file.  Change the text in the arguments, then add your own self- %
007 % contained LaTeX resume, etc.                                         %
008 \documentstyle[12pt,noibmres]{article}
009 \begin{document}
010 \name{George B. Leeman, Jr.}
011 \preferredtitle0{Mr. Leeman}
012 \school0{The University of Michigan}
013 \email{leeman at um.cc.umich.edu}
014 \discipline0{mathematics}
015 \highestdegree0{PhD}
016 \degreeyear0{1972}
017 \oneaddressandphone{
018 2904 Washtenaw Avenue, Apt. 1B\\
019 Ypsilanti, MI 48198\\
020 (313) 434-1815
021 }
022 \immigrationstatus{US citizen}
023 \typeofpositiondesired{Permanent}
024 % Put your own LaTeX commands and text after this line.
    .
    .
    .
    \end{document}

38 OBTAINING FILES FROM SOFTWARE.WATSON.IBM.COM

Appendix A. Obtaining files from software.watson.ibm.com

User input lines are numbered, system responses are unnumbered.

01 ftp internet-gw Access internet gateway via FTP.

VM TCP/IP FTP V2R2
Connecting to INTERNET-GW 129.34.128.85, port 21
220-4.3 BSD UNIX (NRT) #1: Thu Nov 14 19:58:31 EST 1991
220 internet-gw FTP GATEWAY (Version 4.24 Mon Dec 18 17:59:45 EST 1989) ready.
USER (identify yourself to the host):

02 leeman Type userid.

>>>USER leeman
331 Password required for leeman.
Password:

03 Type password, which is not displayed.

>>>PASS ********
200 Specify Remote Destination with SITE command
Command:
04 site software.watson.ibm.com  Specify desired server.
>>>SITE software.watson.ibm.com
220 software.watson.ibm.com FTP server (IBM RT) ready.
Command:

05 user anonymous Log in with userid “anonymous”

>>>USER anonymous
331-Software on this FTP server is furnished by the terms
331-and conditions put forth in the README.FTP file in the
331-pub directory.  By continuing, you agree to get
331-the above mentioned file and agree to these conditions with
331-respect to the software here.  Otherwise, disconnect now.
331 Guest login ok, send ident as password.
Password:

06 Type email address, e.g. leeman@watson.ibm.com

>>>PASS ********
230 Guest login ok, access restrictions apply.
Command:

07 get /pub/README.FTP First retrieve the README file.

>>>PORT 129,34,128,247,190,215
200 PORT command successful.
>>>RETR /pub/README.FTP
150 Opening data connection for /pub/README.FTP (129.34.139.1,1279) (3376 bytes)
226 Transfer complete.

3449 bytes transferred. Transfer rate 30.52 Kbytes/sec.
Command:

08 cd /pub/ibmres-tex Make /pub/ibmres-tex current directory

>>>CWD /pub/ibmres-tex
250 CWD command successful.
Command:

09 get ibmres.sty Retrieve ibmres.sty file.

>>>PORT 129,34,128,247,190,224
200 PORT command successful.
>>>RETR ibmres.sty
150 Opening data connection for ibmres.sty (129.34.139.1,1280) (14851 bytes).
226 Transfer complete.
15164 bytes transferred. Transfer rate 57.88 Kbytes/sec.
Command:

10 get simpres.ltx Retrieve simpres.ltx file.

>>>PORT 129,34,128,247,190,228
200 PORT command successful.
>>>RETR simpres.ltx
150 Opening data connection for simpres.ltx (129.34.139.1,1281) (1959 bytes).
226 Transfer complete.
2026 bytes transferred. Transfer rate 36.84 Kbytes/sec.
Command:

11 get res_u_f.ltx Retrieve res_u_f.ltx file.

>>>PORT 129,34,128,247,190,233
200 PORT command successful.
>>>RETR res_u_f.ltx
150 Opening data connection for res\_u\_f.ltx (129.34.139.1,1282) (5646ytes).
226 Transfer complete.
5805 bytes transferred. Transfer rate 52.30 Kbytes/sec.
Command:

12 quit Sever connection with the server.

>>>QUIT
221 Goodbye.

40 LISTING OF /PUB/IBMRES-TEX/IBMRES.STY FILE.

Appendix B. Listing of /pub/ibmres-tex/ibmres.sty file.

001 %                                                             10/08/93 %
002 % IBM resume style file, /pub/ibmres-tex/ibmres.sty                     %
003 % George B. Leeman, Jr., leeman@watson.ibm.com.                         %
004 % Please read /pub/README.FTP on software.watson.ibm.com before using   %
005 % this file.  Documentation for using this file is in                   %
006 % /pub/ibmres-tex/ibmres.ltx.                                          %
007 %=======================================================================%
008 % The next 140 lines are the RESUME.STY file of Stephen Gildea, with    %
009 % enhancements by Jerry Leichter and George B. Leeman, Jr.              %
010  % This style option is designed to work with the article document
011  % style of LaTeX version 2.09.  Use
012  % \documentstyle[resume]{article}
013
014  % Copyright (c) 1987 by Stephen Gildea
015  % Permission to copy all or part of this work is granted, provided
016  % that the copies are not made or distributed for resale, and that
017  % the copyright notice and this notice are retained.
018  %
019  % THIS WORK IS PROVIDED ON AN "AS IS" BASIS.  THE AUTHOR PROVIDES NO
020  % WARRANTY WHATSOEVER, EITHER EXPRESS OR IMPLIED, REGARDING THE WORK,
021  % INCLUDING WARRANTIES WITH RESPECT TO ITS MERCHANTABILITY OR FITNESS
022  % FOR ANY PARTICULAR PURPOSE.
023
024  % If you make any improvements, I'd like to hear about them.
025
026  % Stephen Gildea
027  % MIT Earth Resources Lab
028  % Cambridge, Mass. 02139
029  % mit-erl!gildea
030  % gildea@erl.mit.edu
031
032  % 23 May 87 version 1
033  % 08 Jun 88 center single \address lines
034  % 02 Sep 88 enhancement to \employer from Jerry Leichter
035  %          LEICHTER-JERRY@cs.yale.edu or <LEICHTER@Venus.YCC.Yale.Edu>
036  % 08 Mar 89 allow \\ in \sectiontitle to make multi-line titles
037
038  \topmargin 0pt
039  \headheight 0pt
040  \headsep 0pt
041  \textheight 9.0in

042  \pagestyle{empty}
043  \parindent 0pt
044  \parskip \baselineskip
045  \topmargin 0in
046  \oddsidemargin -0.1in % was 0in
047  \evensidemargin 0in
048  \textwidth 6.55in % was 6.5in
049
050  \setcounter{secnumdepth}{0}
051
052  % use this to print your name at the top of the page
053  \def\name#1{{\large\bf\centering#1\\[2\baselineskip]}\ignorespaces}
054
055  % one address
056  \def\address#1{{\centering #1\par}}
057
058  % two addresses (say, home and work).  They get printed side by side
059 \def\addresses#1#2{\hbox to \hsize{\@tablebox{#1}\hfil\@tablebox{#2}}}
060
061  % CC: [t] stands for align on top row; {@{}l@{}} indicates
062  % that there is one entry that should be aligned to the left and
063  % surrounded by two blank spaces.
064  % Another example is in ieeeconf.sty, where we have
065  % \def\author#1#2{\begin{tabular}[t]{c@{\extracolsep{8em}}c}#1&#2\end{tabular}}
066  \def\@tablebox#1{\begin{tabular}[t]{@{}l@{}}#1\end{tabular}}
067
068  % CC: \newbox reserves a register of type box for  \@title (p. 244)
069  \newbox\@title
070
071  % Most of the resume goes inside a llist environment.  Within it, use
072  % the \sectiontitle{title} command to begin new sections.  The title
073  % of the section is put in the left margin.  The section text begins
074  % on the next line if the sectiontitle is too long.
075  %
076  % CC: (1) \begin{list}{default_label}{\labelwidth \leftmargin \labelsep
077  % \rightmargin \topsep \parsep \parskip \itemsep \parsep} \item_list
078  % \end{list} are latex commands (see pp.112--115 and p.165--166)
079  % CC: (2) \newenvironment{command_name}[args]{begin_definition}{end_definition}
080  % Modification by George B. Leeman: next line to change left margin width, then
081  % change width accordingly in llist environment.  Also add a multiple
082  % line section title (mlsectiontitle) which allows multiple lines
083  % inside items, formatting them to the proper width.
084  \newlength\leftmarginwidth     \setlength\leftmarginwidth{1 in}
085  \newenvironment{llist}{%

086 %\begin{list}{}{\setlength\labelwidth{1in}%
087  \begin{list}{}{\setlength\labelwidth{\the\leftmarginwidth}%
088   \setlength\leftmargin{\labelwidth}
089   \addtolength\leftmargin{\labelsep}
090   \itemsep 5pt plus 2pt minus 2pt
091   \parsep 10pt plus 2pt minus 2pt
092  %
093  % Set the depth of the title to 0 in case more than one line.
094  % CC: \copy<register> produces a copy of box register <register>.
095  % This command is useful when you want to retrieve the contents of a box
096  % but don't want to destroy the contents.
097  \def\sectiontitle##1{\setbox\@title=\hbox{{\bf\@tablebox{##1}}}\dp\@title=0pt
098    \item[\copy\@title]%
099  % If the title takes more lines than the body, you lose.
100  % \ifdim\wd\@title>\labelwidth \leavevmode \\* \else \fi}
101  % CC: Because for sectiontitiles with \wd\@title>\labelwidth, we always
102  % have \item following them, we do not need an extra line
103   \ifdim\wd\@title>\labelwidth \leavevmode \else \fi} %end of sectiontitle
104 % \def\mlsectiontitle##1{\setbox\@title=\hbox{\@tablebox{##1}}\dp\@title=0 pt
105 %   \item[\copy\@title]}
106   \def\mlsectiontitle##1{\setbox\@title=\hbox{\@tablebox{
107     \parbox[t]{\leftmarginwidth}{\raggedright##1}}}\dp\@title=0 pt
108     \item[\copy\@title]}
109  \def\makelabel##1{##1\hfill} } }{\end{list}}
110
111  % like itemize but bullets are at the left margin
112  \newenvironment{items}{%
113   \begin{list}{$\bullet$}{\setlength\labelwidth{.5em}\setlength\leftmargin{1em}
114  % CC: new formats added by me :-)
115    \parsep 0pt \parskip 0pt \topsep 1pt %\partopsep -2pt
116    \def\makelabel##1{##1\hfill}}}{\end{list}}
117
118  % Use the \employer and \location commands to print the names and
119  % cities of your employers.
120  % \employer always starts a new paragraph, and should be followed by
121  % \location, because no \par after \employer
122  %
123  % \employer* is like \employer but is used for additional information
124  % about the employer.  It does NOT start a new paragraph, must
125  % NOT be followed by \location, and always forces a new line after itself.
126  \def\employer{\@ifstar{\@semployer}{\@employer}}
127  %\def\@employer#1{\par{\sc #1}}
128  \def\@employer#1{{\sc #1}}
129  \def\@semployer#1{{\sc #1}\\}

130  % CC: used when the next environment is \begin{items} \end{items}
131  \def\lastemployer#1{{\sc #1}}
132
133  % The location is always flush right.  It is moved to the next
134  % line if there is not room left on this one.
135  % See the TeXbook Chapter 14.
136  % CC: \penalty<number> produces a penality item. A penality of 10000
137  % discourages (prevents) a break, while a penality of -10000 forces
138  % a break. \finalhyphendemerits<number>: increasing <number> has
139  % the effect of discouraging TEX from ending the nect ot the last line
140  % with a hyphen.
141  \def\location#1{{\unskip\nobreak\hfill\penalty50\hskip2em
142    \hbox{}\nobreak\hfill \hbox{#1}\finalhyphendemerits=0 \\}}
143
144  % This prints the date flush left.
145  % The major shortcoming of this file:
146  % There needs to be an option to print the date in the left margin,
147  % because many people like to do it that way.
148 % \def\dates#1{#1\\[2pt]}
149 % End of RESUME.STY file.                                              %
150 %=======================================================================%
151 % Additions by George B. Leeman, Jr., October 22, 1992.                 %
152 %                                                                      %
153 % New ifs, boxes, dimens, lengths, tokens.                              %
154 \newif\if@nocoverletter
155 \newbox\@onebox
156 \newbox\@oneline
157 \newdimen\@closingwidth
158 \newdimen\@iwidth
159 \newdimen\@longindentation
160 \newdimen\@remainder
161 \newlength{\@fwidth}
162 \newlength{\@oneaddresssize}
163 \newlength{\@twoaddresssize}
164 \newtoks\@tempone
165 \newtoks\@temptwo
166 % Macro definitions, in alphabetical order.                             %
167 \def\abstracts#1{\newpage\pagestyle{empty}%
168   \centerline{{\bf Selected Abstracts}}\vskip 5 ex}
169 \long\def\@append#1\to#2{\@tempone=\expandafter{#1}\@temptwo=\expandafter{#2}%
170   \edef#2{\the\@temptwo\the\@tempone}}
171 \def\availabledate#1{\hbox to \hsize{\@tablebox{{\bf Present Status: }%
172   \@presentstatustext\\{\bf Type of Position Desired: }%
173   \@typeofpositiondesiredtext}\hfil

174   \@tablebox{{\bf Immigration Status: }\@immigrationstatustext\\%
175     {\bf Available Date: }#1}}\ignorespaces}
176 \def\cc#1{\if@nocoverletter\relax\else\@docc{#1}\fi\ignorespaces}
177 \def\closing#1{\if@nocoverletter\relax\else\@setclosing#1\0\fi}
178 \def\coverletter#1{\end{llist}\@ifnull{#1}\@nocoverlettertrue
179   \else\def\@returnaddress{{\ignorespaces#1}}\@makeuniv#1\\\0\fi}
180 \def\degree#1{\def\@degreetext{\ignorespaces#1}}
181 \def\degreeyear0#1{\ignorespaces}
182 \def\discipline0#1{\ignorespaces}
183 \def\@docc#1{\@ifnull{#1}\relax\else\par\noindent\parbox[t]{\textwidth}%
184   {\@hangfrom{\rm cc: }\ignorespaces #1\strut}\par\fi}
185 \def\@doencl#1{\@ifnull{#1}\relax\else\par\noindent\parbox[t]{\textwidth}%
186   {\@hangfrom{\rm encl: }\ignorespaces #1\strut}\par\fi}
187 \def\education{\sectiontitle{Education}}
188 \def\email#1{\def\emailtext{}\@ifnull{#1}\relax\else\def\emailtext{\\#1}\fi
189   \ignorespaces}
190 \def\employerspecificdata#1{\@ifnull{#1}\relax
191   \else{\hbox to \hsize{\@tablebox{{\bf \@employerspecifictext: }#1}}}\fi
192   \ignorespaces}
193 \def\employerspecifictitle#1{\@ifnull{#1}\relax
194   \else\def\@employerspecifictext{\ignorespaces#1}\fi\ignorespaces}
195 \def\employment{\sectiontitle{Employment}}
196 \def\encl#1{\if@nocoverletter\relax\else\@doencl{#1}\fi}
197 \def\gpa#1{\@ifnull{#1}\def\@gpatext{ {\em no GPA.}}%
198   \else\def\@gpatext{ {\em GPA: }\ignorespaces#1}\fi}
199 \def\highestdegree0#1{\ignorespaces}
200 \def\homeaddressandphone#1{\@newsize\@twoaddresssize{\bf Home Address}\\#1\\\0%
201   \def\saveaddress{}\@setlineend\saveaddress{\bf Home Address}\\#1\\\0%
202   \addresses{\saveworkaddress}%
203     {\parbox[t]{\@fwidth}{\raggedright\saveaddress}}%
204   \vskip 4 ex\ignorespaces}
205 \def\honors#1{\sectiontitle{Honors}#1\par}
206 \def\@ifnull#1{\def\temp{#1}\ifx\temp\empty}
207 \def\immigrationstatus#1{\def\@immigrationstatustext{\ignorespaces#1}%
208   \ignorespaces}
209 \long\def\letterbody#1{\if@nocoverletter\relax\else#1\par\fi}
210 \def\letterlabel{\if@nocoverletter\ignorespaces
211   \else\newpage\thispagestyle{empty}\hbox{\vrule\vbox{\hrule\kern6truept
212       \hbox{\kern6truept\vbox to 2truein{\hsize=\@longindentation
213           {\smallheadfont\baselineskip9truept\@savename\\\@returnaddress\par}
214           \vfill\moveright 2truein\vbox{\raggedright\hsize=\@longindentation
215             \everypar{\hangindent 1 em}\rm\@receiveraddress}\vfill}%
216         \kern6truept}\kern6truept\hrule}\vrule}
217   \setcounter{page}{0}\newpage\fi}

218 \def\@makeletterhead{\setcounter{page}{0}\newpage
219   \setlength{\textwidth}{6.25 truein}\hsize=6.25 truein
220   \def\ps@empty{\let\@mkboth\@gobbletwo\def\@oddhead{}\def\@oddfoot{}%
221     \def\@evenhead{}\def\@evenfoot{}\headsep 0 pt\headheight 0 pt
222     \textheight 9 in}%
223   \parindent=0 pt\parskip=0 pt\headsep 6 ex\headheight 12 pt
224   \addtolength{\textheight}{-6 ex}\addtolength{\textheight}{-12 pt}\raggedbottom
225   \interlinepenalty=1000 \pagestyle{myheadings}%
226   \markright{{\rm To: \@savereceiver\hfill\today\hfill Page }}%
227   \@longindentation=4 truein\@remainder=2.25 truein\thispagestyle{empty}
228   \font\smallheadfont=cmr8 at 8 truept\font\largeheadfont=cmr10 at 14.4 truept
229   \hbox{\largeheadfont\@saveuniversity\hfil}\hbox{\smallheadfont\@secondline}%
230   \vskip 2 truept\hrule\vskip 4 truept\hspace*{\@longindentation}%
231   \parbox{\@remainder}{\baselineskip 9 truept\smallheadfont\@sendingaddress\par}
232   \par\vspace{\bigskipamount}\vspace{\bigskipamount}\hspace*{\@longindentation}%
233   \rm\today\par\vspace{\bigskipamount}\@receiveraddress\par
234   \vspace{\bigskipamount}}
235 \def\@makename#1\\#2\0{\def\@savename{{\ignorespaces#1}}}
236 \def\@makereceiver#1\\#2\0{\def\@savereceiver{{\ignorespaces#1}}}
237 \def\@makeuniv#1\\#2\0{\def\@saveuniversity{{\ignorespaces#1}}\@setsecond#2\\\0}
238 \def\miscellaneous#1{\sectiontitle{Miscellaneous}#1\par}
239 \def\@newsize#1#2\\#3\0{\@iwidth=#1\advance\@iwidth by -1 em
240   \global\setbox\@onebox
241     \vbox{\hsize=#1\parshape=2 0 pt #1 0 pt\the\@iwidth\raggedright#2\par}%
242   \loop
243     \global\setbox\@onebox
244       \vbox{\unvbox\@onebox\global\setbox\@oneline\lastbox\unskip\unpenalty}%
245     \global\setbox\@oneline\hbox{\unhbox\@oneline\unskip}%
246     \global\@iwidth=\wd\@oneline
247     \ifdim\ht\@onebox>0 pt
248       \advance\@iwidth by 1 em
249       \ifdim\@fwidth<\the\@iwidth\global\@fwidth=\the\@iwidth\fi
250     \repeat
251   \ifdim\@fwidth<\the\@iwidth\global\@fwidth=\the\@iwidth\fi
252   \@ifnull{#3}\relax\else\@newsize\@twoaddresssize#3\0\fi}
253 \def\objective#1{\begin{llist}\sectiontitle{Objective}#1\par}
254 \def\oneaddressandphone#1{\@newsize\@oneaddresssize#1\emailtext\\\0%
255   \def\saveaddress{}\@setlineend\saveaddress#1\emailtext\\\0%
256   \def\addressbox{\raggedright\saveaddress}%
257   \address{\parbox[t]{\@fwidth}{\addressbox}}\vskip 5 ex}
258 \def\outof#1{\@ifnull{#1}\def\@outoftext{}
259   \else\def\@outoftext{ out of \ignorespaces#1.}\fi\@degreetext\@gpatext
260   \@outoftext\par}
261 \def\place{\employer}

262 \def\preferredtitle0#1{\ignorespaces}
263 \def\presentstatus#1{\def\@presentstatustext{\ignorespaces#1}\ignorespaces}
264 \def\ps#1{\if@nocoverletter\relax\else\ignorespaces#1\par\@startbreaks\fi
265   \ignorespaces}
266 \def\publications#1{\sectiontitle{Publications}#1\par}
267 \def\@receiveraddress{}
268 \def\recipient#1{\if@nocoverletter\relax
269   \else\def\@receiveraddress{\ignorespaces#1}\@makereceiver#1\\\0\@makeletterhead
270   \fi}
271 \def\referenceemail#1{\@reftext\@ifnull{#1}\else{ \em Email: }#1\fi\par}
272 \def\referencename#1{\def\@reftext{\ignorespaces#1}}
273 \def\references{\sectiontitle{References}}
274 \def\@returnaddress{}
275 \def\rightlines#1{\if@nocoverletter\relax\else\def\@sendingaddress
276   {\ignorespaces#1}\@makename#1\\\0\fi}
277 \def\@savename{}
278 \def\@savereceiver{}
279 \def\@saveuniversity{}
280 \def\school0#1{\ignorespaces}
281 \def\@secondline{}
282 \def\@sendingaddress{}
283 \long\def\@setclosing#1\\#2\0{\par\nobreak\bigskip\@stopbreaks\noindent
284   \hspace*{\@longindentation}\@closingwidth=\textwidth
285   \advance\@closingwidth -\@longindentation
286   \parbox{\@closingwidth}{\raggedright\ignorespaces #1\\[6\medskipamount]
287     #2\strut}\par}
288 \def\@setcommas#1\\#2\0{\@ifnull{#1}\relax
289   \else\@append{, {\ignorespaces #1}}\to\@secondline\@setcommas#2\0\fi}
290 \long\def\@setlineend#1#2\\#3\0{%
291   \@append{{\ignorespaces#2}\hangindent=1 em \par}\to#1%
292   \@ifnull{#3}\relax\else\@setlineend#1#3\0\fi}
293 \def\@setsecond#1\\#2\0{\@ifnull{#1}\relax
294   \else\def\@secondline{\ignorespaces #1}\@setcommas#2\\\0\fi}
295 \def\skills#1{\sectiontitle{Skills}#1\par}
296 \def\@startbreaks{\let\\=\@normalcr\interlinepenalty 200
297   \def\par{\@@par\penalty 200 \relax}}
298 \def\@stopbreaks{\interlinepenalty \@M \def\par{\@@par\nobreak}\let\\=\@nobreakcr
299   \let\vspace\@nobreakvspace}
300 \long\def\text#1{\ignorespaces#1\par}
301 \def\typeofpositiondesired#1{\def\@typeofpositiondesiredtext{\ignorespaces#1}%
302   \ignorespaces}
303 \def\when#1{\mlsectiontitle{#1}}
304 \def\workaddressandphone#1{\def\workaddress{}%
305 % \@setlineend\workaddress{\bf Work Address}\\#1\emailtext\\\0\noindent

REFERENCES 47

306   \@setlineend\workaddress{\bf Work Address}\\#1\emailtext\\\0%
307   \def\saveworkaddress{{\parbox[t]{\@twoaddresssize}{\raggedright\workaddress}}}%
308   \ignorespaces}
309 % Initialize various dimensions.
310 \setlength{\@oneaddresssize}{\textwidth}
311 \addtolength{\@oneaddresssize}{-2 in}
312 \setlength{\@twoaddresssize}{\textwidth}
313 \addtolength{\@twoaddresssize}{-1 in}
314 \addtolength{\@twoaddresssize}{-0.5 \@twoaddresssize}
315 \setlength{\@fwidth}{-1 pt}
316 \settowidth{\leftmarginwidth}{\bf Miscellaneous}
317 \addtolength{\leftmarginwidth}{1 em}

References

[1] Stephen Gildea, RESUME.STY, May 23, 1987, with improvements on:

June 8, 1988; September 2, 1988 by Jerry Leichter, March 8, 1989; and October 23, 1992 by George B. Leeman, Jr.

[2] Donald E. Knuth. The TEXbook. Addison-Wesley, Reading, Massachusetts, 1984.

[3] Leslie Lamport. A Document Preparation System L^{A}TEX User’s Guide & Reference Manual Addison-Wesley, Reading, Massachusetts, 1986.